# LC-LLM: Explainable Lane-Change Intention and Trajectory Predictions with Large Language Models

### v1

@@ III-C Fine-tuning
In our work, we utilize an open-source foundational language model, Llama-2-7b-chat [8], as our pre-trained LLM. To achieve parameter-efficient finetuning, we adopt the Low-Rank Adaptation (LoRA) [37] strategy, which freezes the pre-trained model weights and injects trainable rank decomposition matrices into each layer of the Transformer architecture. Additionally, we customize the LLM for our specific prediction task using supervised fine-tuning techniques. The raw dataset utilized in our work originates from the highD dataset [38], which captures human naturalistic vehicle trajectories on German highways. Following the fine-tuning instructions provided for Llama-2, each data sample is formatted to include an input prompt and a corresponding answer, separated by a special token. This formatting is illustrated in the Fig. 4. The answer for each sample is derived from the ground truth obtained from the highD dataset, encompassing future trajectories and intentions for programmatic automatic labeling. In summary, we process highD raw dataset with natural language, format each sample into the Llama format, and then feed it into LLM. Finally, we fine-tune the LLM by aligning the LLM’s output $\{\mathcal{I},\mathcal{T}\}$ with the corresponding ground truth labels $\{\mathcal{I}^{*},\mathcal{T}^{*}\}$. This fine-tuning process is executed through the language modeling loss $\mathcal{L}$ as defined in Equation 4. During fine-tuning, we mask the loss for tokens within the input prompts, focusing backpropagation only on the tokens comprising the answers, in alignment with the fine-tuning approach of Llama-2. By following this approach, our fine-tuned model, LC-LLM, can predict human driving behaviors, including lane change intentions and future trajectories. This capability is crucial for ensuring the safety of ego vehicles within an autonomous driving system.
III-D Interpretability of Prediction
@@
A prevalent limitation in contemporary autonomous driving prediction models lies in their constrained interpretability. This is attributed to the fact that these models generate predictions about the future behaviors or trajectories of the target vehicle through black-box neural networks, offering little explanation behind their prediction results. In our research, we address this limitation by incorporating explanatory requirements into the input prompts in the inference stage, where the weights of our fine-tuned model are fixed. An example of these input prompts during the inference stage is illustrated in Fig. 5. Benefiting from the self-explanation [12, 13] capabilities of LLMs, our fine-tuned model, LC-LLM, predicts lane change intentions and future trajectories and also provides explanations for its prediction results in the inference stage, thereby enhancing the interpretability of the prediction results. This approach facilitates a more transparent and understandable prediction process, which is crucial for the practical application of autonomous driving systems. Some visual examples of the interpretability of our LC-LLM model are shown in Fig. 8.
@@ IV Experimental Results and Analysis
Load_in_bits 4
@@ IV-A Dataset Processing
IV-B Experimental Setup
We utilize the training dataset extracted from the highD dataset to train our proposed LC-LLM model, as well as the reimplemented baseline models: an LSTM-based model[22] and a Transformer-based model[3]. All the models are implemented using the PyTorch framework[39]. Additionally, we employ the DeepSpeed [40] library for distributed training of our LC-LLM model. The training process is conducted on eight A800 GPUs and takes 8 hours for the 7B model. The hyperparameters used for model training are detailed in Table I. To ensure the stability of the experimental results, each experiment is repeated five times, and the average value is reported as the final result.
@@ IV-B Experimental Setup
Figure 7: Prompt ablation studies. The ‘Base’ refers to the model results obtained using only the target vehicle information for prompting. The ‘+Map info’, ‘+SV info’, and ‘+CoT’ indicate the model results when map information, surrounding vehicle information, and CoT statements are sequentially incorporated, respectively.
@@ IV-D Overall Performance
Figure 8: Some visual examples demonstrating the interpretability of our LC-LLM model. In each example, the image above depicts the driving scenario, and the text below presents the corresponding output of our model, which includes the prediction results and their explanations.
@@ IV-D1 Intention Prediction Results and Analysis
Table II shows the overall performance of the proposed LC-LLM model as well as baseline models in intention prediction with four different advanced prediction time T. We can note that our proposed LC-LLM significantly outperforms baseline models for all T, especially for the longer T. For example, when $\textit{T}\in(2,3]$, the average F1 score of LC-LLM-7B is 2.1% higher than that of the Transformer model and 3.6% higher than that of the LSTM. Subsequently, when considering $\textit{T}\in(3,4]$, the average F1 score of LC-LLM-7B surpasses that of the Transformer model by 1.3% and outperforms the LSTM model by 6.3%. Furthermore, in comparison to the LC-LLM-7B model, the LC-LLM-13B model demonstrates improvements across all advanced prediction times. However, upon scaling the model parameters of LC-LLM to 70B, a decline in intention prediction performance for the longer T is observed. The decrease in performance may be due to not having enough data for fine-tuning. The scaling laws[41] suggest that as the number of parameters in a model increases, it should be accompanied by a corresponding increase in the amount of training data for optimal performance. In this case, the LC-LLM with 70B parameters might not have had sufficient fine-tuning data, leading to suboptimal results as the model couldn’t fully utilize its expanded capacity.
@@
Only Intention 0.861
Only Trajectory 0.542 2.032
Intention+Trajectory 0.886 0.478 1.198
@@
Zero-shot(Llama-7B-chat) 0.218 457 2.531 53.795 2928
Fine-tune(LC-LLM-7B) 0.886 0 0.478 1.198 0
IV-D2 Trajectory Prediction Results and Analysis
We assessed the performance of our LC-LLM models in the trajectory prediction task and compared them with baseline models. The results are depicted in Fig. 6. The left panel illustrates the trends of lateral RMSE for trajectory points across different models and prediction horizons, while the right panel shows the trends of longitudinal RMSE. The results indicate a notable improvement in both lateral and longitudinal RMSE for our LC-LLM models in comparison to the baselines. Notably, with larger prediction horizons, our model demonstrates increasing advantages over the baselines in longitudinal RMSE, indicating better robustness concerning long-term prediction horizons. Furthermore, the observed trend suggests that the performance of our LC-LLM model improves as its number of parameters grows, which is particularly evident in the right panel. The relatively small improvement in the left panel is mainly attributed to the fact that the lateral vehicle deviation is minimal in the highway driving scenario, limiting the model’s ability to demonstrate performance differences.
The results of lane change intention prediction task and trajectory prediction task demonstrate that our LC-LLM outperforms all baseline models in long-term (4 s) prediction, achieving a 3.1% improvement in lane change intention prediction, a 19.4% improvement in lateral trajectory prediction, and a 38.1% improvement in longitudinal trajectory prediction, respectively.
IV-E Ablation Study
IV-E1 Prompt Ablation Study
We explored the impact of incorporating contextual prompts on the intention prediction task and the trajectory prediction task, and the results are illustrated in Fig. 7. A series of ablation experiments were conducted, progressively introducing different sources of information—Target Vehicle Information, Map Information, Surrounding Vehicle Information, and Chain-of-Thought statements. The comprehensive model, integrating all contextual prompts, exhibited superior performance compared to the base model with only Target Vehicle Information, achieving an impressive 19.08% improvement in F1 score, a substantial 22.65% reduction in lateral RMSE, and a remarkable 67.85% decrease in longitudinal RMSE. Specifically, the addition of Map Information significantly reduces the lateral error in trajectory prediction, while the inclusion of Surrounding Vehicle Information proved crucial for both intention prediction and future trajectory forecasting. Furthermore, the incorporation of Chain-of-Thought statements yielded notable enhancements in both lane change intention prediction and trajectory prediction. These findings suggest the importance of considering diverse contextual information in prompt engineering with LLM.
IV-E2 Multi-task Ablation Study
IV-F Interpretability Study
In order to demonstrate the interpretability of our LC-LLM, we visualize several sample scenarios alongside the corresponding outputs from our model, as depicted in Fig. 8. The visualization illustrates that our LC-LLM is capable of forecasting lane change intentions and future trajectories of the target vehicle while also providing explanations for the prediction results. Additionally, the analysis of our LC-LLM’s output indicates that our model not only focuses on the state information of the target vehicle itself but also considers the interactive information of surrounding vehicles to predict the target vehicle’s lane change intention and future trajectory. For example, as shown in the second case in the figure, the target vehicle exhibits a low velocity with no vehicles ahead and no apparent interaction with adjacent vehicles. The explanation given by our LC-LLM is consistent with this driving scenario. In the third case, our LC-LLM elucidates that given the high velocity of the left-front vehicle and the absence of other vehicles on the left, the target vehicle has an opportunity to execute a left lane change. Moreover, in the fourth case, with a slow-moving truck positioned ahead, our LC-LLM predicts that the target vehicle may make a left lane change to avoid collision with the truck. The explanations provided by our model for these scenarios serve as evidence that our LC-LLM possesses a deep understanding of the driving scenario and makes reasonable predictions for the target vehicle. In contrast to previous methodologies that solely generate prediction results, our LC-LLM model not only produces prediction results but also provides corresponding explanations, thereby exhibiting better interpretability than previous methods.
@@ V Conclusion
In this paper, we introduce LC-LLM, an explainable lane change prediction model that not only forecasts lane change intentions and trajectories but also provides explanations for its predictions. We reconceptualize the task of lane change prediction as a language modeling problem and employ a supervised fine-tuning technique to fine-tune the LLM specifically for this task. In this way, we successfully leverage the strong common sense reasoning capabilities and self-explanation abilities of the LLM to address the challenge of lane change prediction. Our extensive experiments conducted on highD dataset show that our LC-LLM improves the accuracy of predicting lane change intention and trajectory as well as significantly augments the interpretability of the prediction results.
Currently, this paper only verifies the effectiveness of our approach in highway scenarios. In the future, our approach should be extended to a wider range of traffic scenarios, and additional experiments on urban scene datasets should be conducted. Furthermore, our LC-LLM exhibit longer inference time compared to baseline models, making the compression of LLMs through knowledge distillation an important area for future research. Moreover, the proposed approach adopts a target-centric strategy to make predictions for one interested vehicle, which results in considerable computational demands when there are multiple interested vehicles in the same scene. Therefore, developing a method that can simultaneously forecast lane change intentions and trajectories for multiple vehicles is an essential direction for future work.

### v2

@@ III-D Fine-tuning
In our work, we utilize an open-source foundational language model, Llama-2-13b-chat [9], as our pre-trained LLM. To achieve parameter-efficient finetuning, we adopt the Low-Rank Adaptation (LoRA) [40] strategy, which freezes the pre-trained model weights and injects trainable rank decomposition matrices into each layer of the Transformer architecture. Additionally, we customize the LLM for our specific prediction task using supervised fine-tuning techniques. The raw dataset utilized in our work originates from the highD dataset [41], which captures human naturalistic vehicle trajectories on German highways. Following the fine-tuning instructions provided for Llama-2, each data sample is formatted to include an input prompt and a corresponding answer, separated by a special token. This formatting is illustrated in the Fig. 5. The answer for each sample is derived from the ground truth obtained from the highD dataset, encompassing programmatically automatically labeled future trajectories, driving intentions and CoT reasoning. In summary, we process highD raw dataset with natural language, format each sample into the Llama format, and then feed it into LLM. Finally, we fine-tune the LLM by aligning the LLM’s output $\{\mathcal{C},\mathcal{I},\mathcal{T}\}$ with the corresponding ground truth labels $\{\mathcal{C}^{*},\mathcal{I}^{*},\mathcal{T}^{*}\}$. This fine-tuning process is executed through the language modeling loss $\mathcal{L}$ as defined in Equation 4. During fine-tuning, we mask the loss for tokens within the input prompts, focusing backpropagation only on the tokens comprising the answers, in alignment with the fine-tuning approach of Llama-2. By following this approach, our fine-tuned model, LC-LLM, can predict human driving behaviors and corresponding CoT reasoning. This capability is crucial for ensuring the safety of ego vehicles within an autonomous driving system.
III-E Interpretability of Prediction
@@
A prevalent limitation in contemporary autonomous driving prediction models lies in their constrained interpretability. This is attributed to the fact that these models generate predictions about the future behaviors or trajectories of the target vehicle through black-box neural networks, offering little explanation behind their prediction results. In our research, we address this limitation in two distinct ways. First, we fine-tune the CoT reasoning capabilities of our LC-LLM. By doing so, our model learns the domain knowledge and rules in driving, which enhances transparency and reliability of lane change predictions, leading to more accurate and explainable predictions. Second, we incorporate explanatory requirements into the input prompts in the inference stage, where the weights of our fine-tuned model are fixed. Benefiting from the self-explanation [10, 11] capabilities of LLMs, our fine-tuned model, LC-LLM, not only predicts lane change intentions and future trajectories but also provides corresponding CoT reasoning and explanations for its prediction results, thereby enhancing the interpretability of the prediction results. This approach facilitates a more transparent and understandable prediction process, which is crucial for the practical application of autonomous driving systems. Some visual examples of the interpretability of our LC-LLM model are shown in Fig. 8.
@@ IV Experimental Results and Analysis
Warmup_steps 600
Load_in_bits 8
@@ IV-A Dataset Processing
IV-B Experimental Setup
We utilize the training dataset extracted from the highD dataset to train our proposed LC-LLM model, as well as the reimplemented baseline models: an LSTM-based model[3] and a Transformer-based model[4]. All the models are implemented using the PyTorch framework[42]. Additionally, we employ the DeepSpeed [43] library for distributed training of our LC-LLM model. The training process is conducted on eight A800 GPUs and takes 11 hours for the 13B model. The hyperparameters used for model training are detailed in Table I. To ensure the stability of the experimental results, each experiment is repeated five times, and the average value is reported as the final result.
@@ IV-D1 Intention Prediction Results and Analysis
Table II shows the overall performance of the proposed LC-LLM model as well as baseline models in intention prediction with four different advanced prediction time T. We can note that our proposed LC-LLM significantly outperforms baseline models for all T, especially for the longer T. For example, when $\textit{T}\in(2,3]$, the average F1 score of LC-LLM is 13.8% higher than that of the Transformer model and 15.5% higher than that of the LSTM. Additionally, when $\textit{T}\in(3,4]$, the average F1 score of LC-LLM surpasses that of the Transformer model by 17.7% and outperforms the LSTM model by 25.7%. These results demonstrate that our LC-LLM performs well even at the early stage of a lane change maneuver when lateral displacement is minimal.
IV-D2 Trajectory Prediction Results and Analysis
We assessed the performance of our LC-LLM models in the trajectory prediction task and compared them with baseline models. The results are depicted in Fig. 6. The left panel illustrates the trends of lateral RMSE for trajectory points across different models and prediction horizons, while the right panel shows the trends of longitudinal RMSE. The results indicate a notable improvement in both lateral and longitudinal RMSE for our LC-LLM models in comparison to the baselines. Notably, with larger prediction horizons, our model demonstrates increasing advantages over the baselines in lateral RMSE and longitudinal RMSE, indicating better robustness concerning long-term prediction horizons.
The results of lane change intention prediction task and trajectory prediction task demonstrate that our LC-LLM outperforms all baseline models in long-term (4 s) prediction, achieving a 17.7% improvement in lane change intention prediction, a 64.4% improvement in lateral trajectory prediction, and a 66.1% improvement in longitudinal trajectory prediction, respectively.
Figure 7: Components ablation studies. The ‘Base’ refers to the model results obtained using only the target vehicle information for prompting. The ‘+Map info’, ‘+SV info’, and ‘+Thought’ indicate the model results when map information, surrounding vehicle information, and CoT reasoning are sequentially incorporated, respectively.
IV-E Ablation Study
IV-E1 Components Ablation Studies
We explored the impact of incorporating different components on the trajectory prediction task, and the results are illustrated in Fig. 7. A series of ablation experiments were conducted, progressively introducing different components—Target Vehicle Information, Map Information, Surrounding Vehicle Information, and Chain-of-Thought reasoning. The comprehensive model, integrating all contextual prompts and CoT reasoning component, exhibited superior performance compared to the base model with only Target Vehicle Information, achieving a substantial 10.64% reduction in lateral RMSE, and a remarkable 29.57% decrease in longitudinal RMSE. Specifically, the addition of Surrounding Vehicle Information proved crucial for the future trajectory forecasting. Furthermore, the incorporation of CoT reasoning yielded notable enhancements in trajectory prediction.
@@ IV-E1 Components Ablation Studies
Only Intention 0.948
Only Trajectory 0.257 0.798
Intention+Trajectory 0.971 0.210 0.655
IV-E2 Multi-task Ablation Study
Figure 8: Some visual examples demonstrating the interpretability of our LC-LLM model. In each example, the image on the left depicts the driving scenario, and the text on the right presents the corresponding output of our model, which includes the prediction results and their reasoning and explanations.
IV-F Interpretability Study
IV-F1 Visualization of Results
In order to demonstrate the interpretability of our LC-LLM, we visualize several sample scenarios alongside the corresponding outputs from our model, as depicted in Fig. 8. The visualization illustrates that our LC-LLM is capable of forecasting lane change intentions and future trajectories of the target vehicle while also providing corresponding CoT reasoning and explanations for the prediction results. For example, in the first case in the figure, with a slow-moving truck positioned ahead, the target vehicle need to overtaking to gain some speed advantage. And given the high velocity of the left front vehicle while the low velocity of right front vehicle, the target vehicle has an opportunity to execute a left lane change. The explanation given by our LC-LLM is consistent with this driving scenario. These explanations provided by our model for these scenarios serve as evidence that our LC-LLM possesses a deep understanding of the driving scenario and makes reasonable and reliable predictions for the target vehicle. In contrast to previous methodologies that solely generate prediction results, our LC-LLM model exhibit better interpretability.
IV-F2 Quantitative Evaluation
To quantitatively evaluate the interpretability of our LC-LLM, we evaluate the accuracy of CoT reasoning. We randomly selected 100 samples from the test dataset and manually compared the accuracy of the CoT reasoning output by our LC-LLM model with the ground truth in the dataset, scoring them accordingly. The scoring rule is that 10 points will be deducted for each error, omission, or addition of a notable feature, and 50 points will be deducted for a potential behavior error. Our quantitative evaluation result is that our model scores 97.2. This result demonstrates that our LC-LLM model excels in generating accurate and interpretable predictions, confirming the effectiveness of our methods in enhancing the transparency and reliability of autonomous driving systems.
@@ IV-F2 Quantitative Evaluation
Zero-shot(Llama-13B-chat) 0.210 4 1.405 63.366 4
Fine-tune(LC-LLM) 0.971 0 0.210 0.655 0
@@ IV-G Zero-shot Study
IV-H Robustness Evaluations
Figure 9: Some visual examples evaluating the robustness of our LC-LLM. These test samples are from outside the distribution of the highD dataset.
As mentioned in [44], safety-critical events are important to test the model’s robustness. In this experiment, we evaluated the robustness of our LC-LLM by creating four safety-critical scenarios not included in the HighD datasets. The first scenario involves a left lane change with the left front vehicle emergency braking, with speeds of 0 to 50 km/h in 10 km/h increments and relative distances of 10 to 100 meters in 10-meter increments, resulting in 60 samples. The second scenario tests a right lane change with the right front vehicle emergency braking, using the same speed and distance ranges, also resulting in 60 samples. The third scenario involves a left lane change with the left rear vehicle accelerating, with speeds of 100 to 150 km/h in 10 km/h increments and the same distance range, resulting in 60 samples. The fourth scenario examines a right lane change with the right rear vehicle accelerating, using the same speed and distance ranges as the third scenario, producing 60 samples.
The results, depicted in Fig. 9, highlight the robustness of our LC-LLM across these four challenging scenarios. In scenario (a), where the right front vehicle stops as the target vehicle changes lanes to the right, the LC-LLM accurately predicted the necessary adjustments to avoid a collision. Scenario (b) shows the right rear vehicle accelerating as the target vehicle changes lanes to the right; here, the LC-LLM successfully anticipated the increased speed of the rear vehicle and adjusted the trajectory accordingly. In scenario (c), the model handled the left front vehicle’s deceleration effectively during the target vehicle’s left lane change, maintaining safe distances and smooth transitions. Finally, in scenario (d), the LC-LLM demonstrated its ability to manage the left rear vehicle’s acceleration during the target vehicle’s left lane change, ensuring a safe maneuver. These results affirm the model’s capability to handle out-of-distribution cases, enhancing its applicability in real-world autonomous driving scenarios.
IV-I Limitations
Our LC-LLM, while demonstrating significant improvements in lane change intention and trajectory prediction accuracy and interpretability, has limitations. Firstly, our model has been tested exclusively on the HighD dataset, which primarily comprises highway scenarios. To ensure the robustness and generalizability of our model, it is essential to evaluate its performance on more diverse datasets such as NuScenes [45] and Waymo [46], which encompass a wider variety of complex urban and suburban traffic scenarios. While incorporating detailed map information from these complex urban scenarios into input prompts presents a considerable challenge. Consequently, the development of an appropriate prompt that accurately describe high-definition maps remains an objective for our future work. Secondly, our LC-LLM exhibits slower inference speed compared to baseline models. Specifically, the inference time for LC-LLM-13B is 26.72 seconds with a batch size of 32, significantly longer than the 0.011 seconds and 0.014 seconds for the LSTM and Transformer models, respectively. This slower inference time is partly due to the larger model parameters, posing a challenge for real-time applications. Future work should focus on optimizing the model for faster inference and exploring techniques such as knowledge distillation to reduce the computational load.
@@ V Conclusion
In this paper, we introduce LC-LLM, an explainable lane change prediction model that not only forecasts lane change intentions and trajectories but also provides CoT reasoning and explanations for its predictions. We reconceptualize the task of lane change prediction as a language modeling problem and employ a supervised fine-tuning technique to fine-tune the LLM specifically for this task. In this way, we successfully leverage the strong common sense reasoning capabilities and self-explanation abilities of the LLM to address the challenge of lane change prediction. Our extensive experiments conducted on highD dataset show that our LC-LLM improves the accuracy of predicting lane change intention and trajectory as well as significantly augments the interpretability of the prediction results. Future work includes extending our approach to urban driving scenarios, reducing inference times through knowledge distillation to improve model speed and efficiency, and developing methods to predict lane change intentions and trajectories for multiple vehicles simultaneously.